# Towards the understanding of heavy quarks hadronization: from leptonic to heavy-ion collisions

### v2

@@ IV Prospects and future developments
In this section, we discuss prospects and goals for the future, also touching some topics and open points connected to heavy-flavour hadronization that could not be deepened in the previous sections. These includes recent measurements that have not yet conducted to firm conclusions, either because still preliminary or because affected by large uncertainties, but that by exploring different observables are pointing to new effects and directions of search for the future. We also discuss new ideas and measurements, which have the potential to lead toward an understanding of the comparison between experimental measurements performed by different collaborations and the mismatch with some theoretical calculations.
@@
The comparison of the measurements performed by ALICE and LHCb at mid and forward rapidity, respectively, show tantalising signs of a possible rapidity dependence of the baryon-to-meson ratio in pp, p–Pb, and Pb–Pb collisions [276, 219, 180, 218, 78, 274, 72, 220] for both the $\Lambda_{\mathrm{c}}^{+}$ and $\Xi_{\mathrm{c}}^{+,0}$ baryons, as displayed in Fig. 27. However, the current measurements do not allow for conclusions, either because of the large uncertainties (including those on the BR in the $\Xi_{\mathrm{c}}^{+,0}$ case) or because of inconsistencies in the accessed $p_{\mathrm{T}}$ intervals or in the definition and ranges used for multiplicity and centrality. It is envisioned that the LHC collaborations in the future will, as far as possible, perform measurements in the same centrality intervals and transverse momentum, facilitating easier comparisons and reducing potential ambiguities.
@@
There are several areas of interest in extending the current modelling of string hadronization. Some have already been mentioned above, including modelling of exotica and the heavy-ion limit of strings. Further areas of interest include considering colour reconnections in the initial state, the collective behaviour of strings, interactions between final state hadrons, studies into baryon production mechanisms, and even looking beyond Lund strings. Here we elaborate on just a few of these areas of interest.
@@
As extensively discussed in Ref. [297], the $z_{\mathrm{||}}^{\mathrm{ch}}$ distribution of D-meson tagged jets measured by the ALICE Collaboration is quite well described by Monte Carlo event generators. Recently, ALICE reported a first measurement of the same distribution for $\Lambda_{\mathrm{c}}^{+}$-tagged jets. The comparison with D-meson tagged jet data done with the same jet and hadron kinematic selection is reported in Fig. 28. Though the uncertainties prevent firm conclusions, the data suggest that $\Lambda_{\mathrm{c}}^{+}$ carry a smaller fraction of jet momentum along the jet axis with respect to D mesons. This trend contrasts with that expected by the default version of PYTHIA 8 (Monash tune), while it agrees with that from PYTHIA 8 CR-BLC. The trend expected by Monash reflects the tendency in jet fragmentation that heavier particles carry a larger fraction of jet momentum, which in a collinear fragmentation picture can be related to energy conservation. The data therefore suggest that either the $\Lambda_{\mathrm{c}}^{+}$ production is favoured in jets with higher momentum and softer fragmentation, either intrinsically or because modified by the interplay with the underlying event, or that the $\Lambda_{\mathrm{c}}^{+}$ momentum vector is less aligned with the jet one, either as an intrinsic feature of charm-jet fragmentation to $\Lambda_{\mathrm{c}}^{+}$ baryons (not expected however in Monash) or because of a angular decorrelation at hadronization time. A partly related study is that of the azimuthal correlation of $\Lambda_{\mathrm{c}}^{+}$ baryons with charged particles produced in the event.
A first preliminary measurement, performed by ALICE [305], indicates that the yields of low-$p_{\mathrm{T}}$ associated particles are large, both in the near-side and in the opposite direction to the $\Lambda_{\mathrm{c}}^{+}$ (away-side jet region), and resemble those observed only at much higher D-meson $p_{\mathrm{T}}$ ($p_{\mathrm{T}}\gtrsim 8$ $\mathrm{GeV}/c$) in correlations of D mesons with charged particles. Naively one could expect that, except for the initial-parton energy correlation, the fragmentation (parton-shower) and hadronization of the near- and away-side jets should proceed almost independently. It is thus surprising that the larger near-side yield, which is consistent with the softer jet fragmentation suggested by the $z_{\mathrm{||}}^{\mathrm{ch}}$ analysis, is accompanied by a difference in the away-side jet.
@@
The unique combination of high rates, wide acceptance, strong particle identification capabilities, and excellent secondary vertex determination of the proposed ALICE 3 detector, also provide a toolbox for the search for light nuclei with charm. At LHC energies, the most promising candidates are the c-deuteron, c-triton and c-3He. The existence of such weakly decaying bound states with lifetimes similar to those of other unbound charmed baryons is debated in the literature. In case these states are bound and under the assumption that their abundance in collisions of nuclei is described by the SHM, the yields are large enough to bring their experimental discovery within reach in the next decades and to additionally test different hadronization models.
Exotica
@@
The $\chi_{c1}(3872)$ is an exotic particle that was first observed by the Belle Collaboration [315], and then confirmed and studied by other experiments at electron-positron [316] and hadron colliders [317, 318, 319, 320, 321]. The nature of this particle is still not fully understood and interpretations in terms of conventional charmonium (a bound state of charm-anticharm quarks), D-meson molecule, tetraquark, or an admixture these states have been proposed [322]. Heavy-ion collisions, as well as high multiplicity pp collisions, offer a unique window into the properties of this poorly understood hadron. In these systems, promptly produced $\chi_{c1}(3872)$ hadrons can interact with other particles in the nucleus and/or those produced in the collision. The influence of these interactions on the observed $\chi_{c1}(3872)$ yields provides information that can help discriminating between the various models of its structure, as well as give insight into the dynamics of the bulk of the particles produced in these collisions. The LHCb Collaboration has measured a significant decrease in the ratio of prompt $\chi_{c1}(3872)$ to $\psi$(2S) cross sections in pp collisions with increasing charged-particle multiplicity [323]. These data were interpreted in terms of breakup of the $\chi_{c1}(3872)$ hadron due to interactions with comoving particles produced in the event, for both compact and molecular models of $\chi_{c1}(3872)$ structure [324, 325]. A relatively-compact tetraquark is preferred in this case, but it is currently discussed how conclusive this comparison is [326, 327, 328, 329]. The LHCb Collaboration has also recently performed the same measurement in p–Pb collisions [330]. In this collision system the ratio is measured to be higher than the one measured in pp collisions. It must be noted that the evolution of $\psi$(2S) production with particle multiplicity is not fully understood yet. This complicates the interpretation of $\chi_{c1}(3872)$ data, when $\psi$(2S) is used as a reference. The CMS collaboration has measured the $\mathrm{\sigma}_{\chi_{c1}(3872)}/\sigma_{\psi(2S)}$ ratio in Pb–Pb collisions at high $p_{\mathrm{T}}$, and found that the ratio is further enhanced relative to pp and p–Pb collisions, although the uncertainties and the significantly different $p_{\mathrm{T}}$ intervals of the measurements preclude drawing firm conclusions [331]. Calculations based on quark coalescence show that production rates of $\chi_{c1}(3872)$ hadrons in AA collisions are sensitive to its structure. In these models, the production of compact tetraquarks is expected to be enhanced over hadronic molecules [332, 333]. However, a recent transport calculation reaches the opposite conclusion [326]. Late-stage interactions in the hadron gas phase of a heavy-ion collision can also modify the yields [334]. It is currently unknown where the crossover between the suppressing effect of breakup and the enhancing effect of coalescence may occur. The increase of the ratio of cross sections $\mathrm{\sigma}_{\chi_{c1}(3872)}/\sigma_{\psi(2S)}$ from pp to p–Pb to Pb–Pb collisions may indicate that the exotic $\chi_{c1}(3872)$ hadron experiences different dynamics in the medium than those relevant for a “conventional” charmonium state like $\psi$(2S). If the experiments could in the future cover similar $p_{\mathrm{T}}$ intervals in the different collision systems and multiplicities/centralities, possibly better assessing the lower $p_{\mathrm{T}}$ region, these measurements will provide new constraints to determine the internal $\chi_{c1}(3872)$ structure as well as to model parton transport and hadronization in nuclear collisions. In this context, it will also be important to precisely measure the production cross section of other exotic states, like the $\mathrm{T}^{+}_{cc}$.
@@
A related possible future implementation in PYTHIA is the inclusion of exotica. Not only does QCD-CR allow beyond LC topologies such as junctions, but it also allows for more complicated structures called junction networks. These junction networks are string configurations that consist of multiple junctions within the structure and provide a natural way within the string picture to represent tetraquarks, pentaquarks, etc. For example, the simplest of these structures are junction-antijunction configurations which consist of two colour-charged endpoints and two anticolour-charged endpoints as shown in Fig. 29. When these junction-antijunction configurations are encountered as high mass systems, PYTHIA currently handles these structures by making breaks on the string piece connecting the junction to the antijunction, and then fragmenting each junction system separately. However, should a junction-antijunction system have a low invariant mass, one could propose that the system does not fragment and instead can be treated as a tetraquark. The same argument holds for larger junction networks. The inclusion of exotica is not yet implemented in PYTHIA and it would not be straightforward to implement the required significant modifications to the existing code, however junction networks are already present and could provide a sensible starting point for modelling such structures.

### v3

@@ IV Prospects and future developments
In this section, we outline the prospects and objectives for future research, addressing additional topics and unresolved issues related to heavy-flavour hadronization that were not fully explored in earlier sections. This includes recent measurements that have yet to yield definitive conclusions, either due to their preliminary nature or because they are affected by significant uncertainties. However, the examination of various observables hints at new effects and potentials for future investigation. We also present new ideas and experimental approaches that may help resolve discrepancies between measurements conducted by different collaborations and their divergence from certain theoretical predictions.
@@
The comparison of the measurements performed by ALICE and LHCb at mid and forward rapidity, respectively, show tantalising signs of a possible rapidity dependence of the baryon-to-meson ratio in pp, p–Pb, and Pb–Pb collisions [289, 27, 21, 26, 16, 287, 28, 25] for both the $\Lambda_{\mathrm{c}}^{+}$ and $\Xi_{\mathrm{c}}^{+,0}$ baryons, as displayed in Fig. 27. However, the current measurements do not allow for conclusions, either because of the large uncertainties (including those on the BR in the $\Xi_{\mathrm{c}}^{+,0}$ case) or because of inconsistencies in the accessed $p_{\mathrm{T}}$ intervals or in the definition and ranges used for multiplicity and centrality. As mentioned in Sect. III.5, also the rapidity dependence of the evolution with multiplicity of the $\mathrm{D_{s}^{+}}/\mathrm{D^{0}}$ ratio in p–Pb collisions is not yet experimentally established. It is envisioned that the LHC collaborations in the future will, as far as possible, perform measurements in the same centrality intervals and transverse momentum, facilitating easier comparisons and reducing potential ambiguities.
@@
There are several areas of interest in extending the current modelling of string hadronization. Some have already been mentioned above, including modelling of exotic hadrons and the heavy-ion limit of strings. Further areas of interest include considering colour reconnections in the initial state, the collective behaviour of strings, interactions between final state hadrons, studies into baryon production mechanisms, and even looking beyond Lund strings. Here we elaborate on just a few of these areas of interest.
@@
As extensively discussed in Ref. [309], the $z_{\mathrm{||}}^{\mathrm{ch}}$ distribution of D-meson tagged jets measured by the ALICE Collaboration is quite well described by Monte Carlo event generators. Recently, ALICE reported a first measurement of the same distribution for $\Lambda_{\mathrm{c}}^{+}$-tagged jets. The comparison with D-meson tagged jet data done with the same jet and hadron kinematic selection is reported in Fig. 28. Though the uncertainties prevent conclusions, the data may suggest that $\Lambda_{\mathrm{c}}^{+}$ baryons carry a smaller fraction of jet momentum along the jet axis with respect to D mesons. This would contrast with what expected by the default version of PYTHIA 8 (Monash tune), while it would agree with what PYTHIA 8 CR-BLC expects. The trend expected by Monash reflects the tendency in jet fragmentation that heavier particles carry a larger fraction of jet momentum, which in a collinear fragmentation picture can be related to energy conservation. The data therefore may suggest that either the $\Lambda_{\mathrm{c}}^{+}$ production is favoured in jets with higher momentum and softer fragmentation, either intrinsically or because modified by the interplay with the underlying event, or that the $\Lambda_{\mathrm{c}}^{+}$ momentum vector is less aligned with the jet one, either as an intrinsic feature of charm-jet fragmentation to $\Lambda_{\mathrm{c}}^{+}$ baryons (not expected however in Monash) or because of a angular decorrelation at hadronization time. A partly related study is that of the azimuthal correlation of $\Lambda_{\mathrm{c}}^{+}$ baryons with charged particles produced in the event.
A first preliminary measurement, performed by ALICE [317], indicates that the yields of low-$p_{\mathrm{T}}$ associated particles are large, both in the near-side and in the opposite direction to the $\Lambda_{\mathrm{c}}^{+}$ (away-side jet region), and resemble those observed only at much higher D-meson $p_{\mathrm{T}}$ ($p_{\mathrm{T}}\gtrsim 8$ $\mathrm{GeV}/c$) in correlations of D mesons with charged particles. Naively, one could expect that, except for the initial-parton energy correlation, the fragmentation (parton-shower) and hadronization of the near- and away-side jets should proceed almost independently. It is thus surprising that the larger near-side yield, which is consistent with the softer jet fragmentation suggested by the $z_{\mathrm{||}}^{\mathrm{ch}}$ analysis, is accompanied by a difference in the away-side jet.
@@
The unique combination of high rates, wide acceptance, strong particle identification capabilities, and excellent secondary vertex determination of the proposed ALICE 3 detector, also provide a toolbox for the search for light nuclei with charm. At LHC energies, the most promising candidates are the c-deuteron, c-triton and c-3He. The existence of such weakly decaying bound states with lifetimes similar to those of other unbound charmed baryons is debated in the literature. In case these states are bound and under the assumption that their abundance in collisions of nuclei is described by the SHM [58, 327, 328, 329, 330], the yields are large enough to bring their experimental discovery within reach in the next decades and to additionally test different hadronization models.
Exotic hadrons
@@
The $\chi_{c1}(3872)$ is an exotic particle that was first observed by the Belle Collaboration [331], and then confirmed and studied by other experiments at electron-positron [332] and hadron colliders [333, 334, 335, 336, 337]. The nature of this particle is still not fully understood and interpretations in terms of conventional charmonium (a bound state of charm-anticharm quarks), D-meson molecule, tetraquark, or an admixture these states have been proposed [338]. Heavy-ion collisions, as well as high multiplicity pp collisions, offer a unique window into the properties of this poorly understood hadron. In these systems, promptly produced $\chi_{c1}(3872)$ hadrons can interact with other particles in the nucleus and/or those produced in the collision. The influence of these interactions on the observed $\chi_{c1}(3872)$ yields provides information that can help discriminate between the various models of its structure, as well as give insight into the dynamics of the bulk of the particles produced in these collisions. The LHCb Collaboration has measured a significant decrease in the ratio of prompt $\chi_{c1}(3872)$ to $\psi$(2S) cross sections in pp collisions with increasing charged-particle multiplicity [339]. These data were interpreted in terms of breakup of the $\chi_{c1}(3872)$ hadron due to interactions with comoving particles produced in the event, for both compact and molecular models of $\chi_{c1}(3872)$ structure [340, 341]. A relatively-compact tetraquark is preferred in this case, but it is currently discussed how conclusive this comparison is [342, 343, 344, 345]. The LHCb Collaboration has also recently performed the same measurement in p–Pb collisions [346]. In this collision system the ratio is measured to be higher than the one measured in pp collisions. It must be noted that the evolution of $\psi$(2S) production with particle multiplicity is not fully understood yet. This complicates the interpretation of $\chi_{c1}(3872)$ data, when $\psi$(2S) is used as a reference. The CMS collaboration has measured the $\mathrm{\sigma}_{\chi_{c1}(3872)}/\sigma_{\psi(2S)}$ ratio in Pb–Pb collisions at high $p_{\mathrm{T}}$, and found that the ratio is further enhanced relative to pp and p–Pb collisions, although the uncertainties and the significantly different $p_{\mathrm{T}}$ intervals of the measurements preclude drawing firm conclusions [347]. Calculations based on quark coalescence show that production rates of $\chi_{c1}(3872)$ hadrons in AA collisions are sensitive to its structure. In these models, the production of compact tetraquarks is expected to be enhanced over hadronic molecules [348, 349]. However, a recent transport calculation reaches the opposite conclusion [342]. Late-stage interactions in the hadron gas phase of a heavy-ion collision can also modify the yields [350]. It is currently unknown where the crossover between the suppressing effect of breakup and the enhancing effect of coalescence may occur. The increase of the ratio of cross sections $\mathrm{\sigma}_{\chi_{c1}(3872)}/\sigma_{\psi(2S)}$ from pp to p–Pb to Pb–Pb collisions may indicate that the exotic $\chi_{c1}(3872)$ hadron experiences different dynamics in the medium than those relevant for a “conventional” charmonium state like $\psi$(2S). If the experiments could in the future cover similar $p_{\mathrm{T}}$ intervals in the different collision systems and multiplicities/centralities, possibly better assessing the lower $p_{\mathrm{T}}$ region, these measurements will provide new constraints to determine the internal $\chi_{c1}(3872)$ structure as well as to model parton transport and hadronization in nuclear collisions. In this context, it will also be important to precisely measure the production cross section of other exotic states, like the $\mathrm{T}^{+}_{cc}$.
@@
A related possible future implementation in PYTHIA is the inclusion of exotic hadrons. Not only does QCD-CR allow beyond LC topologies such as junctions, but it also allows for more complicated structures called junction networks. These junction networks are string configurations that consist of multiple junctions within the structure and provide a natural way within the string picture to represent tetraquarks, pentaquarks, etc. For example, the simplest of these structures are junction-antijunction configurations which consist of two colour-charged endpoints and two anticolour-charged endpoints as shown in Fig. 29. When these junction-antijunction configurations are encountered as high mass systems, PYTHIA currently handles these structures by making breaks on the string piece connecting the junction to the antijunction, and then fragmenting each junction system separately. However, should a junction-antijunction system have a low invariant mass, one could propose that the system does not fragment and instead can be treated as a tetraquark. The same argument holds for larger junction networks. The inclusion of exotic hadrons is not yet implemented in PYTHIA and it would not be straightforward to implement the required significant modifications to the existing code, however junction networks are already present and could provide a sensible starting point for modelling such structures.
@@
The points above outline an exciting program of future measurements and theoretical developments that will require significant effort from both the experimental and theoretical communities involved in the topic. Understanding hadronization may also shed light on the properties and degrees of freedom of the system in the partonic phase before and during the formation of hadrons. From the experimental side, some of the mentioned measurements may not be within reach for another 10-15 years, i.e., until the completion of the ALICE and LHCb detector upgrades proposed for Run 5 at the LHC. Nonetheless, achieving a satisfactory understanding and modelling of such a fundamental process of nature is certainly worth the effort.